OpenStudy (anonymous):

the hypotenuse of a right triangle is 26 millimeters. One leg of the right triangle is 10 millimeters. What is the length of the other leg? 18 millimeters 20 millimeters 24 millimeters 28 millimeters

OpenStudy (anonymous):

\[a^2 +b^2 = h^2 \] a,b= length of legs h = hypotenuse

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Put a= 10 and h = 26, then find b